OpenStudy (anonymous):

(-b+b)(2+7) when b=5

OpenStudy (kidrah69):

\[(-5+5)(2+7)\] solve

OpenStudy (anonymous):

I dont know help plz medal

OpenStudy (kidrah69):

-5+5=?

OpenStudy (anonymous):

use a calculater if you have to and if you are aloud

OpenStudy (anonymous):

0 lol

OpenStudy (kidrah69):

right now whats 7+2

OpenStudy (anonymous):

im not allowd

OpenStudy (kidrah69):

7+2=?

OpenStudy (pinklion23):

kidrah69 is explaining this very simply

OpenStudy (anonymous):

9

OpenStudy (kidrah69):

9 x 0 = ?

OpenStudy (anonymous):

0

OpenStudy (pinklion23):

-5+5-0 so )*9=? your answer you can do it :)

OpenStudy (pinklion23):

ta da

OpenStudy (anonymous):

thank uuuuuu

OpenStudy (kidrah69):

?

OpenStudy (pinklion23):

@kidrah69 you helped him very good great job @Rafyy have any more :)

OpenStudy (kidrah69):

\[9\times0=?\]

OpenStudy (pinklion23):

he got it he said 0

OpenStudy (anonymous):

yeah later I think lol

OpenStudy (kidrah69):

good job and thanks @pinklion23 <3